OpenStudy (anonymous):

Factor completely 64x2 – 1

OpenStudy (whpalmer4):

This is a difference of two squares, an important form to recognize when factoring. \[(a-b)(a+b) = a^2 + ab - ab - b^2 = a^2 -b^2\] So, what values of a and b give you your equation?